OpenStudy (anonymous):

Could someone help me with (-45x^-8x^6y^4z^0/9x^2y^-2z)^-2 PLEASE!

OpenStudy (anonymous):

ooh man... u need the simplified version of this?

OpenStudy (anonymous):

I can't tell what's being multiplied and what's an exponent.

OpenStudy (anonymous):

Does this help?

OpenStudy (anonymous):

Yes. First thing to do is work within the parenthesis. The numbers are easy: -45/9 gives us -5. For exponents, we add them if multiplying, subtract them if dividing. So with x, -8 + -6 - 2 = -16. Likewise for y and z. At this point you should have (-5*x^(-16)*y^(6)*z^(-1))^(-2). Now you can distribute the ^(-2) across each term: (-5)^(-2) = 1/25 (or 1/(5^2)) x^(-16*-2) = x^32 y^(6*-2) = y^(-12) (or 1/(y^(12)) z^(-1*-2) = z^2 So you should end up with: ((x^32)*(z^2))/(25(y^12)). Of course, you should also double check this to make sure I did it right in my head and to make sure you understand the process.

OpenStudy (anonymous):

Grr....docx